OpenStudy (anonymous):

help me factor 6x^2+29x-5

OpenStudy (anonymous):

(6x-1)(x+5)

Parth (parthkohli):

Hint:\[6 \cdot (-5) = -30\\30 \cdot (-1) = 29 \ \]

OpenStudy (anonymous):

\[6x^2 + 29x - 5\] You could do what base said or use the quadratic formula.

OpenStudy (anonymous):

=6x^+30x-x-5= 6x(x+5) - 1(x+5) = (6x-1) (x+5) Ans.

Parth (parthkohli):

@base421 Also, Manita, who has asked the question, seeks help. Thank you.